How does (2,0); (3,0) equial 3x+2y=6
can you explain so I can help my daugther

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

(2,0) is the x intercept, where the line crosses the x axis.
(0,3) is the y intercept, where the line crosses the y axis.
3x+2y=6
3(2)+2(0)=6
2 times 0 = 0 so it goes away
---------------
3(0)+2(3)=6
3 times 0 = 0 so it goes away

Answer by solver91311(24713)



In the first place, if you can't spell "daughter" correctly, you probably are the daughter. Don't know what you are trying to accomplish with that story.

In the second place, an ordered pair, such as is not "equal" to a two-variable linear equation such as .

Ordered pairs are elements of the solution set of such an equation. In fact there are an infinite number of ordered pairs that satisfy the given equation. In order to determine whether a particular ordered pair is a member of the solution set, substitute the first value in the ordered pair for in the equation and substitute the second value for in the equation. Then do the indicated arithmetic. If you end up with a true statement, in this case something that looks like , then the ordered pair is an element of the solution set. On the other hand if you end up with an absurdity, something like , then you know the ordered pair is NOT an element of the solution set.

John

My calculator said it, I believe it, that settles it